OpenStudy (phagey):

Write an equation in point-slope form for the line through the given point with the given slope. (1, 8); m = -1/6 A. y - 1 = -1/6(x - 8) B. y + 8 = -1/6(x + 1) C. y - 8 = -1/6(x - 1) D. y + 8 = -1/6(x - 1

OpenStudy (michele_laino):

here we have to apply this formula: \(y-y_0=m(x-x_0)\) where \((x_0,y_0)=(1,8)\), and \(m=-1/6\)

OpenStudy (phagey):

is it B @Michele_Laino

OpenStudy (michele_laino):

no, since the left side, has to be equal to \(y-8\)

OpenStudy (phagey):

i really dont know how to do this @Michele_Laino

OpenStudy (phagey):

so its C @Michele_Laino

OpenStudy (michele_laino):

yes! That's right!

OpenStudy (phagey):

can you help me with 1 more @Michele_Laino

OpenStudy (michele_laino):

ok!

OpenStudy (phagey):

The equation of the line that goes through the points (-1, 1) and (3, 9) is y = 2x + 3. A. True. B. False @Michele_Laino

OpenStudy (michele_laino):

the slope \(m\), of the line which passes at those points you provided, is: \[m = \frac{{{y_2} - {y_1}}}{{{x_2} - {x_1}}} = \frac{{9 - 1}}{{3 - \left( { - 1} \right)}} = \frac{{9 - 1}}{{3 + 1}} = ...?\]

OpenStudy (phagey):

so its true @Michele_Laino

OpenStudy (michele_laino):

please wait, try to complete such compèutation first

OpenStudy (michele_laino):

computation*

OpenStudy (michele_laino):

\(m=...?\)

OpenStudy (phagey):

when i graph the y=2x+3 is going through 4 and -2 @michele_laino

OpenStudy (michele_laino):

It is not possible

OpenStudy (michele_laino):

for example, I replace, into the equation \(y=2x+3\), \(x=-1\) at the right side, so I get: \(y=2x+3=2 \cdot (-1)+3=...\)

OpenStudy (michele_laino):

what is \(y\) ?